OpenStudy (anonymous):

A basketball player makes 8 of 12 free throws in the first game of the season. If she shoots with the same accuracy in the second game, how many of the 15 free throws she attempts will she make?

OpenStudy (anonymous):

8/12=x/15, right?

OpenStudy (anonymous):

Yep..!! Ur right

OpenStudy (anonymous):

Multiply 15 to both sides?

OpenStudy (anonymous):

Cross Multiply. x = 8 x 15/12 = 10

OpenStudy (anonymous):

Got it! Thanks!